OpenStudy (anonymous):

There are four categories:- A, B, C, and D... If all of A are B, 15% of D is C, half of C is B, quarter of D is A, and A, compared to the other has the least number. So what is the percentage of C over the total number of A, B, C and D?

OpenStudy (aravindg):

can u wrte the given info mathematically ?

OpenStudy (anonymous):

The question was given in this particular way, N i dun see how to change its form

OpenStudy (anonymous):

any idea???

OpenStudy (anonymous):

the question is worth 6 marks....

hartnn (hartnn):

related to which topic? i tried using venn diagram, but cannot proceed ahead....

OpenStudy (anonymous):

its just logic maths.....

OpenStudy (anonymous):

AND ALSO HOW IS THIS POSSIBLE? quarter of D is A===>25% of D is A and 15% of D is C WHICH MEANS size of c must be smaller than A

OpenStudy (anonymous):

But that is the question.... its for the senior year placement, so they've been having the ame question for years, n i dont think they've had problems with the question

hartnn (hartnn):

only 15 % of D in C, not entire... if D had 100 elements, out of 100, 15 are in C and 25 are in A

OpenStudy (nubeer):

not sure if right or wrong but here is my try Suppose D=D C=0.15D B = (0.15D / 2) = 0.075D A = 0.25D As all of these are percentages hence there total will be equal to 1 so D+0.15D + 0.075D +0.25D = 1 D = 0.67 Percentage of C = 0.15( 0.67) = 0.101 *100 = 10.17%

OpenStudy (anonymous):

@nubeer i dont think there are decimals involved as they hv gvn space enough for 2 digits only....

OpenStudy (nubeer):

i have just converted them from percentage into decimal.. u can multiply each of them by 100 it will turn back to normal..

OpenStudy (anonymous):

wont that be a bit too much

OpenStudy (nubeer):

its upto u.. if u want to convert decimal or making it whole number doesn't make any difference

OpenStudy (anonymous):

but i dun think its correct

OpenStudy (nubeer):

might be but thats my logic can be wrong or right.

OpenStudy (anonymous):

m not blaming u @nubeer ... i noe hw i stared at the question wen i first saw it

OpenStudy (anonymous):

WELL I GOT 15%

OpenStudy (anonymous):

not sure thats right

OpenStudy (nubeer):

just suppose in D category there are 100 people... that might make easy for u to understand. just substitute D by 100..

OpenStudy (anonymous):

@sauravshakya hw d'you get it....maybe if i noe the method i'll be able to see whether its right or wrong??? and hw tough is it??

OpenStudy (anonymous):

WELL I MADE THE SETS THAT SATISFIES YOUR ALL GIVEN CONDITIONS...(NOT A NICE METHOD) A={d,e,f,g,h} B={d,e,f,g,h,u,v,w,x} C={a,b,c,u,v,w} D={a,b,c,d,e,f,g,h,i,j,k,l,m,n,o,p,q,r,s,t}

OpenStudy (anonymous):

BUT, THIS ALSO SATISFIES YOUR ALL GIVEN CONDITIONS A={d,e,f,g,h} B={d,e,f,g,h,u,v,w,x,y,z} C={a,b,c,u,v,w} D={a,b,c,d,e,f,g,h,i,j,k,l,m,n,o,p,q,r,s,t}

OpenStudy (anonymous):

AND THIS GIVES 14.29%

OpenStudy (anonymous):